OpenStudy (summerncaela):

PLEASE HELP I NEED 100% WILL MEDAL!

OpenStudy (3mar):

I need the question, not the medal

OpenStudy (summerncaela):

its 3 and 4

OpenStudy (3mar):

Ok.

OpenStudy (3mar):

You already know the equations of motion!

OpenStudy (summerncaela):

I mean, i've been missing school quite a bit because of sickness so, i don't fully understand how to do it

OpenStudy (3mar):

I ask Allah to heal you and be better than before.

OpenStudy (summerncaela):

well can you help me with this

OpenStudy (3mar):

so who fill in the table?

OpenStudy (summerncaela):

i did

OpenStudy (3mar):

when?

OpenStudy (summerncaela):

today

OpenStudy (3mar):

so can you refer to row 3?

OpenStudy (summerncaela):

i know how to do question one an two but im unclear on how to do 3 and 4

OpenStudy (3mar):

what is the magnitude of g?

OpenStudy (3mar):

Ok I will help

OpenStudy (3mar):

what is g? its value at you?

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (summerncaela):

do you mean 9.8???

OpenStudy (summerncaela):

i'm so lost

OpenStudy (3mar):

yes, as somewhere they put it as 10 m/s2

OpenStudy (3mar):

step by step In Sha' Allah you will get it

OpenStudy (summerncaela):

ok ru going to help me and show me how?

OpenStudy (3mar):

Of course.Will not be late for any help.

OpenStudy (summerncaela):

please, carry on and help me

OpenStudy (3mar):

Take that and see it when I am talking

OpenStudy (3mar):

Are you there?

OpenStudy (3mar):

hey

OpenStudy (summerncaela):

yes

OpenStudy (summerncaela):

what does f stand for?

OpenStudy (3mar):

got it?

OpenStudy (3mar):

F stands for the additional force applied to the rock. I just need its effect to increase the acceleration. ok?

OpenStudy (summerncaela):

im trying to find the instantaneous velocity at 4,5, and 6 seconds,

OpenStudy (3mar):

please be with me one by one to hit the final answer later, ok?

OpenStudy (summerncaela):

@agent0smith can you help???

OpenStudy (3mar):

now (at the third second from falling) what is the accel.?

OpenStudy (summerncaela):

can you just show me the steps on how to do it, like write it down or something???? find t=4 and ill do the other two

OpenStudy (3mar):

you need 3? or 4?

OpenStudy (3mar):

3 or 4?

OpenStudy (summerncaela):

ohh my gosh, for #3 it says the acceleration increased by 10m/s^2 after falling for 3 seconds, calculate the instantaneous velocity at t=4, t=5, t=6 seconds and you find the instantaneous velocity for t=4 and show the stepss

OpenStudy (3mar):

we are now at 3 a=9.8+10=19.8 m/s2 you have calculated the v already and it is 30 m/s then you have the initial velocity + acceleration now solve for the state 4 as follows t=1 (as we have a new value for accle.) for fifth state take t=2 as the previous reason

OpenStudy (danjs):

A change in velocity depends on your acceleration and how long you accelerate for \[\Delta v = a*t\] \[v _{2}=v _{1}+a*t\] a new velocity is equal to the old velocity plus acceleration times time.

OpenStudy (summerncaela):

im looking for instantaneous

OpenStudy (3mar):

instantaneous = velocity you already have calculated.

OpenStudy (danjs):

This says that the acceleration changes from 10 to 20 at t=3 seconds For t=0 to t=3 \[v _{2}=v _{1}+10*t\] for t=3 onwards \[v _{2}=v _{1}+20*t\]

OpenStudy (danjs):

right before the acceleration goes to 20, the velocity at t=3 is v = 0 + 10*3 = 30 m/s

OpenStudy (danjs):

For time t is larger than t=3, you use v = 30 + 20*t

OpenStudy (summerncaela):

what exactly is the formula for instantaneous velocity

OpenStudy (summerncaela):

@DanJS

OpenStudy (3mar):

Dan has already included this is his expressions!

OpenStudy (summerncaela):

ok chill

OpenStudy (danjs):

it is a limit as time goes to zero of change in position x over change in time t \[\large v=\lim_{\Delta t \rightarrow 0}\frac{ \Delta x }{ \Delta t }=\frac{ dx }{ dt }\]

OpenStudy (danjs):

all the velocities used are considered instantaneous at whatever time t. t=1 v=10 t=2 v=20

OpenStudy (danjs):

For a constant acceleration, the following holds Final velocity = initial velocity + acceleration * time \[\large v _{f}=v _{i}+a*t\] If the acceleration changes from 10 to 20 at time t=3, then at t=3 and forwards, the velocity will be what it was at t=3 plus the acceleration 20 times time after t=3

OpenStudy (danjs):

For time t=3, the velocity is v=30, for times after t=3, the velocity is the 30 plus the acceleration times time (t-3) \[v _{f}=30+20*(t-3)\]